OpenStudy (anonymous):

the real and imaginary solution for x^4-32=14x^2

OpenStudy (anonymous):

can someone plz help me solve this

terenzreignz (terenzreignz):

It's important to be familiar with equations which may not be quadratic but are quadratic in FORM. Take this \[\Large x^4 -32 = 14x^2\] Certainly, it is quartic (highest exponent is 4) but notice that if you make the substitution \(\large u = x^2\) \[\Large u^2 - 32 =14u\] Can you try to solve for u?

OpenStudy (anonymous):

give me 5 min

OpenStudy (anonymous):

u has two solution u=16 u=-2

terenzreignz (terenzreignz):

That's good. Now we substitute back... \[\Large u = 16 = x^2\\\Large u = -2 = x^2\] Now, can you find the four possible solutions for x?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

can u plz help

terenzreignz (terenzreignz):

We have u = 16 and u = -2, right? We undo the substitution and switch back, we now have \[\Large x^2 = 16 \\\Large \] and \[\Large x^2 = -2\] Solve both of these...

OpenStudy (anonymous):

but how

terenzreignz (terenzreignz):

How? LOL You solved the more difficult \(\large u^2 - 32 =14 u\) but you can't solve \(\Large x^2 = 16 \) ? :3 How about if we bring 16 to the left side... \(\Large x^2 - 16 = 0\)

OpenStudy (anonymous):

(x+4)(x-4)=0

terenzreignz (terenzreignz):

And...? x = ?

OpenStudy (anonymous):

-4 and 4

terenzreignz (terenzreignz):

Okay, that's two of them... now what about \(\large x^2 = -2\) ? Just get the square root of both sides...

OpenStudy (anonymous):

x+4=0 x=-4

terenzreignz (terenzreignz):

^I'm well aware... now get started on \(\large x^2 = -2 \)

OpenStudy (anonymous):

(x-1)(x+2)=0

OpenStudy (anonymous):

x=1 x=-2

terenzreignz (terenzreignz):

No... try again... I told you, get the square root of both sides... and for your information... \[\Large (x-1)(x+2) = x^2 +x -2\]